# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.5 Proof of Theorem 4
Obviously, we have $E(\boldsymbol{\varphi}_{\mathcal{A}^{c}}\boldsymbol{\varphi}_{\mathcal{A}^{c}}^{\top})=N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}^{c}}$. Define ${\bf Y}_{t,\mathcal{A}}=(y_{j_{1}t},\dots,y_{j_{d}t})$ for any $1\leq t\leq T$. Then, we have $E(\sqrt{NT^{-1}}\sum_{t=1}^{T}{\bf Y}_{t,\mathcal{A}}\sqrt{NT^{-1}}\sum_{t=1}^{T}{\bf Y}_{t,\mathcal{A}}^{\top})=N\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}}$ and $E(\boldsymbol{\varphi}_{\mathcal{A}^{c}}\sqrt{NT^{-1}}\sum_{t=1}^{T}{\bf Y}_{t,\mathcal{A}}^{\top})=N\sqrt{\omega_{T}/T}\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}}=O(N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}})$. Define
$\mathbf{\Sigma}_{U}=\left(\begin{array}[]{cc}N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}^{c}}&N\sqrt{\omega_{T}/T}\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}}\\ N\sqrt{\omega_{T}/T}\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}^{c}}&N\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}}.\end{array}\right)$
@@
$\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}=\boldsymbol{E}^{\top}\boldsymbol{E}+\boldsymbol{F}^{\top}\boldsymbol{F}+2\boldsymbol{E}^{\top}\boldsymbol{F}$
@@
$\mathbb{P}\left(\left|\boldsymbol{F}^{\top}\boldsymbol{F}+2\boldsymbol{E}^{\top}\boldsymbol{F}\right|\geq\epsilon\nu_{N}\right)\leq\frac{3}{N^{t}}\rightarrow 0,$
by $d=o\left(\lambda_{\min}(\mathbf{R})\operatorname{tr}\left(\mathbf{R}^{2}\right)^{1/2}/(\log N)^{C}\right)$, where $\nu_{N}=\left[2\operatorname{tr}\left(\mathbf{R}^{2}\right)\right]^{1/2},t=t_{N}:=C\epsilon/8\cdot v_{N}/\left[\lambda_{\max}(\mathbf{R})\log N\right]\rightarrow\infty,\epsilon_{N}:=(\log N)^{C}/\left[v_{N}\lambda_{\min}(\mathbf{R})\right]\rightarrow 0$. $\Box$

### v2

@@ A.5 Proof of Theorem 4
Obviously, we have $E(\boldsymbol{\varphi}_{\mathcal{A}^{c}}\boldsymbol{\varphi}_{\mathcal{A}^{c}}^{\top})=N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}^{c}}$. Define ${\bf Y}_{t,\mathcal{A}}=(y_{j_{1}t},\dots,y_{j_{d}t})$ for any $1\leq t\leq T$. Then, we have $E(\sqrt{NT^{-1}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf Y}_{t,\mathcal{A}}\sqrt{NT^{-1}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf Y}_{t,\mathcal{A}}^{\top})=NT^{-1}\sum_{t=1}^{T}E(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})^{2}\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}}=O(N\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}})$ and
$E(\boldsymbol{\varphi}_{\mathcal{A}^{c}}\sqrt{NT^{-1}}\sum_{t=1}^{T}{\bf Y}_{t,\mathcal{A}}^{\top})=N\sqrt{T^{-1}}\sum_{t=1}^{T}E\Big{\{}\omega_{T}^{-1/2}h_{t}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\Big{\}}\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}}=O(N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}}).$
Let $\mathbf{\Sigma}_{22}=NT^{-1}\sum_{t=1}^{T}E(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})^{2}\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}}$, $\mathbf{\Sigma}_{21}=N\sqrt{T^{-1}}\sum_{t=1}^{T}E\Big{\{}\omega_{T}^{-1/2}h_{t}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\Big{\}}\mathbf{\Sigma}_{u,\mathcal{A},\mathcal{A}^{c}}$ and $\mathbf{\Sigma}_{12}=N\sqrt{T^{-1}}\sum_{t=1}^{T}E\Big{\{}\omega_{T}^{-1/2}h_{t}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\Big{\}}\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}}.$ Define
$\mathbf{\Sigma}_{U}=\left(\begin{array}[]{cc}N\mathbf{\Sigma}_{u,\mathcal{A}^{c},\mathcal{A}^{c}}&\mathbf{\Sigma}_{12}\\ \mathbf{\Sigma}_{21}&\mathbf{\Sigma}_{22}\end{array}\right).$
@@
$\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}=\boldsymbol{E}^{\top}\boldsymbol{E}+\boldsymbol{F}^{\top}\boldsymbol{F}+2\boldsymbol{E}^{\top}\boldsymbol{F}.$
@@
$P\left(\left|\boldsymbol{F}^{\top}\boldsymbol{F}+2\boldsymbol{E}^{\top}\boldsymbol{F}\right|\geq\epsilon\left[2\operatorname{tr}\left(\mathbf{R}^{2}\right)\right]^{1/2}\right)\leq\frac{3}{N^{t}}\rightarrow 0,$
by $d=o\left(\lambda_{\min}(\mathbf{R})[\operatorname{tr}\left(\mathbf{R}^{2}\right)]^{1/2}/(\log N)^{C}\right)$, where
$t=t_{N}:=C\epsilon/8\cdot\left[2\operatorname{tr}\left(\mathbf{R}^{2}\right)\right]^{1/2}/\left[\lambda_{\max}(\mathbf{R})\log N\right]\rightarrow\infty,$
and $\epsilon_{N}:=(\log N)^{C}/\left(\left[2\operatorname{tr}\left(\mathbf{R}^{2}\right)\right]^{1/2}\lambda_{\min}(\mathbf{R})\right)\rightarrow 0$. $\Box$